(tan^2 x + 1)/(tan x + cot x)
How do you simplify this to a single trig function using other trig identities?

Answer by stanbon(75887)

(tan^2 x + 1)/(tan x + cot x)
How do you simplify this to a single trig function using other trig identities?
----------------------
sec^2/[(sin/cos) + (cos/sin)]
---------------------
(1/cos^2)/[(sin^2+cos^2)/(sin*cos)]
------
= (1/cos^2)/(1/(sin*cos))
------
= sin/cos
------
= tan(x)
======================
Cheers,
Stan H.